OpenStudy (anonymous):

Find the range if y = 3x + 12 for x = {0.4, 0.6, 1.8, 2.2}.

OpenStudy (aroub):

y = 3x + 12 x = {0.4, 0.6, 1.8, 2.2}. y=3(0.4)+12 y= 13.2 <---- R y=3x+12 y=3(0.6)+12 y=13.8 <--- R y=3x+12 y=3(1.8)+12 y= 17.4 <--- R y=3x+12 y=3(2.2)+12 y=18.6 <--- R ( I realllly hate solving with decimals! :(( lol )